# An “opinion reproduction number” for infodemics in a bounded-confidence content-spreading process on networks

### v1

@@ V Quantifying content spread
Now that we have predicted the onset of infodemics in our bounded-confidence content-spreading model on configuration-model networks, we measure the sizes (i.e., the numbers of nodes) in the resulting dissemination trees. The size of a dissemination tree equals the total number of content shares. Studying dissemination trees allows us to explore the long-term behavior of our content-spreading model. These calculations thereby complement our examination of early-stage spreading in section IV.
@@ V.1 Analysis of the total number of content shares
Given the number of nodes and the degree distribution of a configuration-model graph $G$, the content state $x_{0}$, the receptiveness parameter $c$, and the distribution $\phi(x)$ of opinions, we again use generating functions (as in section IV.1) to estimate the total number of content shares (i.e., the size) of a dissemination tree $G_{0}$. This analysis treats content spread as a percolation process. See Newman (2018) for a detailed description of this approach.
Let $S$ be the fraction of a network’s nodes that spread a piece of content. There are thus a total of $SN$ content shares. Node opinions are independent of connectivity in our content-spreading model, so the probability that a node is in the “spreading set” $\mathcal{S}$ equals the probability that it is receptive to the content multiplied by the probability that it belongs to the connected component with the source node. Let $u$ be the mean probability of not being connected to this component via a particular neighbor. We then have
@@
where we recall that $p_{k}$ is the probability a node has degree $k$ and $g_{0}(u)$ is the generating function of the degree distribution of the network.
To calculate $S$, we need to determine $u$. A particular node $i$ not being connected to the spreading set $\mathcal{S}$ via a particular node $j$ can occur either because it is not receptive to the content or because it is not in the same component as the content. In other words, it has been isolated from the content spread because of other non-receptive nodes. For a node $j$ with $k$ edges aside from the one to node $i$, this occurs with probability $1-s(x_{0},c)+s(x_{0},c)u^{k}$, Therefore,
@@
where $q_{k}$ again denotes the probability of having $k$ neighbors other than the spreading edge and $g_{1}(u)$ again denotes the generating function of the excess degree distribution of the network. By inspection, $u=1$ is always a solution; this solution entails not having an infodemic. We are interested in whether or not there are also solutions $u\in(0,1)$.
In some special cases, it is possible to solve for $u$ analytically, but it is typically difficult (or even impossible) to do so even when $g_{1}(u)$ has a simple closed-form expression. However, it is effective to find the root $u^{*}\in(0,1)$ of section V.1 numerically using either an explicit expression or an approximation of the generating function of the excess degree distribution Newman (2018). One then substitutes $u^{*}$ into eq. 10 to obtain an approximation of $S$. The solution $u^{*}$ yields an approximation of $S$ when there is an infodemic.
In fig. 5, we show an example where we compare our approximation of the total number of content shares ($SN$) to numerical simulations in configuration-model networks of three different sizes. For each size (200, 2400, and 4800 nodes), we generate 1000 configuration-model networks with degree sequences from a Poisson distribution with mean $\lambda=5$. We plot histograms of the total number of content shares, and we observe good agreement between our analytical and numerical results.
Figure 5: Histograms of the total number of content shares across 1000 trials for 200-node, 2400-node, and 4800-node configuration-model networks. The dashed vertical line indicates the number of content shares (i.e., the dissemination-tree sizes) that we predict from our analysis. In each realization, we generate an $N$-node configuration-model network with a degree sequence that we obtain from a Poisson distribution with mean $\lambda=5$. We draw the node opinions $x_{i}$ uniformly at random from $(0,1)$, and we set the content state to be $x_{0}=0.5$ and the receptiveness parameter to be $c=0.2$. In each realization, we draw a new degree sequence and a new set of node opinions.
@@ VI Summary statistics for dissemination trees
Figure 6: The effect of varying the network size $N$ on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spread model on configuration-model networks. The solid curves give the means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate an $N$-node configuration-model network with a degree sequence from a Poisson distribution with mean $\lambda=5$. We vary $N$ from $100$ to $5000$ in increments of $100$. Each realization has different initial node opinions, which we draw uniformly at random from $(0,1)$. In each realization, we draw a new degree sequence and a new set of node opinions. The content state is $x_{0}=0.5$ and the receptiveness parameter is $c=0.2$. Both the total number of content shares (i.e., the dissemination-tree size) and the width grow linearly with $N$. The longest-path length grows quickly at first as we increase $N$, and then it grows much more slowly with $N$. The structural virality appears to saturate at a constant value for sufficiently large $N$.
@@
Figure 7: The effect of varying the expected mean degree $\lambda$ of a network on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spread model on configuration-model networks. The solid curves give the means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate a 2000-node configuration-model network with a degree sequence from a Poisson distribution with mean $\lambda$. We vary $\lambda$ from $1$ to $40$ in increments of 1. Each realization has different initial node opinions, which we draw uniformly at random from $(0,1)$. In each realization, we draw a new degree sequence and a new set of node opinions. The content state is $x_{0}=0.5$ and the receptiveness parameter is $c=0.2$. The total number of content shares grows initially and then saturates at about $800$. The width also tends to increase with $\lambda$, with a possible plateau in a small interval near $\lambda=20$. The longest-path length and structural virality increase at first and then decrease, eventually leveling off at a constant value for sufficiently large expected mean degrees. In our simulations, the longest-path length has a maximum at about $\lambda=4$ and structural virality has a maximum at about $\lambda=3$.
@@
Figure 8: The effect of varying the receptiveness parameter $c$ on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spread model on configuration-model networks. We vary $c$ from $0.02$ to $0.48$ in increments of $0.02$. The solid curves give the means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate a 2000-node configuration model network with a degree sequence from a Poisson distribution with mean $\lambda=5$. Each realization has different initial node opinions, which we draw uniformly at random from $(0,1)$. In each realization, we draw a new degree sequence and a new set of node opinions. The content state is $x_{0}=0.5$. The total number of content shares (i.e., the dissemination-tree size) and the width are very small until about $c=0.1$, and then they grow with $c$. Both growth rates seem roughly linear, although the curve for the total number of content shares seems to be concave down and the receptiveness experiences some slower growth with $c$ before settling into linear growth. The longest-path length increases initially with $c$ before seeming to reach a maximum and then decaying. The structural virality also appears to increase initially with $c$ before seeming to reach a maximum and then decaying to a constant value. The standard deviations of the longest-path length and the structural virality are large.
@@
Figure 9: The effect of varying the content state $x_{0}$ on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spread model on configuration-model networks with initial node opinions from a uniform distribution. We vary $x_{0}$ from $0.02$ to $0.98$ in increments of $0.02$. The solid curves show the means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate a 2000-node configuration-model network with a degree sequence from a Poisson distribution with mean $\lambda=5$. Each realization has different initial node opinions, which we draw uniformly at random from $(0,1)$. In each realization, we draw a new degree sequence and a new set of node opinions. When $x_{0}\in(0.2,0.8)$, each of these four summary statistics is roughly constant. This is not surprising because we draw initial node opinions uniformly at random, so the transmission probability for content to spread along an edge depends only on $c$ in this interval (see eq. 4). The symmetry of these summary statistics with respect to $x_{0}$ is also clear from eq. 4. Content states that are closer to the boundary produce dissemination trees with fewer total content shares, smaller widths, and shorter longest paths. However, the structural virality of such “extreme” content states is slightly larger.
@@
Figure 10: The effect of varying the content state $x_{0}$ on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spread model on configuration-model networks with initial node opinions that we draw from a Gaussian distribution with mean $0.5$ and standard deviation $0.1$. We vary $x_{0}$ from $0.02$ to $0.98$ in increments of $0.02$. The solid curves show the means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate a 2000-node configuration-model network with a degree sequence from a Poisson distribution with mean $\lambda=5$.
@@
In this section, we perform numerical experiments and study the output of our bounded-confidence content-spread model on configuration-model networks. We quantify its features by computing four summary statistics of the dissemination trees that we obtain in our simulations.
@@
  • Length of a longest adoption path. An adoption path is a path in a dissemination tree from the source node to another node in the tree Oh and Porter (2018). We measure the “depth” of content spread of the content by calculating the length of a longest adoption path.
  • Width. One can arrange a dissemination tree $G_{0}$ with the source node at the top (i.e., at level 0), nodes with adoption-path length 1 at level 1, nodes with adoption-path length 2 at level 2, and so on. The width of a dissemination tree is the largest number of nodes that adopt the content in a level. Mathematically, the width is
    $\max_{a\in\{1,\ldots,l_{*}\}}\text{(number of nodes with adoption path length }a)\,,$
    where $l_{*}$ is the length of the longest adoption path.
  • Structural virality. Structural virality (i.e., the Wiener index) was used by Goel et al. Goel et al. (2016) to measure the viral nature of content spread. The structural virality $v$ is the mean shortest-path length between nodes in the dissemination tree $G_{0}$. Mathematically, the structural virality is
@@
One natural question to ask is how the number of nodes of a network affects the spread of content on that network. We study this question by increasing the network size $N$ for fixed expected mean degree $\lambda$ (see fig. 6). With this construction, the expected number of edges remains constant as we increase $N$. As implied by our analysis in section V, the total number of content shares grows linearly with $N$ (with a slope indicated by eq. 10). The dissemination-tree width also appears to grow linearly with $N$. The longest-path length and structural virality change more dramatically for small $N$ than for large $N$. The former appears to grow very slowly for large $N$, and the latter appears to eventually saturate at a constant value.
@@
In fig. 7, we examine the impact of increasing the expected mean degree $\lambda$ on the four summary statistics. The total number of content shares increases with $\lambda$ before saturating at a constant value. The width also tends to increase with $\lambda$, with a possible plateau in a small interval near $\lambda=20$. The longest-path length and structural virality have maxima for small values of $\lambda$. When $\lambda$ is large, it is seemingly possible to for the content to be adopted by all receptive and connected nodes in just a few steps, yielding a large dissemination tree that reaches many new nodes in a single time step, short adoption paths, and small structural viralities. The maxima in the longest-path length and structural virality indicate that a spread process persists longer (and hence travels farther from the source node) when nodes have fewer neighbors on average (but still have large enough degree for an infodemic to occur).
@@
Varying the receptiveness parameter $c$ (see fig. 8) yields similar dissemination-tree statistics as varying the expected mean degree. We again observe maxima in the longest-path length and the structural virality, likely due to larger receptiveness values favoring faster content spread, although these statistics have large standard deviations in this experiment. We observe evidence of a phase transition in the total number of content shares at the value of $c$ that we predicted on our analysis in section IV.1. There also appears to be an accompanying phase transition in the dissemination-tree width.
@@
We also examine the impact of the content state on our dissemination-tree statistics for two different initial opinion distributions (see eq. 3). In fig. 9, we show results of simulations in we draw the initial node opinions from a uniform distribution on the interval $(0,1)$. In fig. 10, we show results of simulations in which we draw the initial node opinions on $\mathbb{R}$ from a Gaussian distribution with mean $0.5$ and standard deviation $0.1$. As expected, we observe smaller values of all four summary statistics when the content state lies in the tails of a opinion distribution $\phi(x)$ (or near the boundary of the distribution interval for the uniform distribution). We observe maxima in the total number of content shares and dissemination-tree width at the distribution mean (and also for a very large interval around it for the uniform distribution). For the Gaussian distribution, the content states that yield the largest longest-path lengths and structural viralities occur away from the mean (and are symmetric), indicating that content can spread farther from the source node in these situations then when the content state equals the mean initial opinion. These dissemination trees are longer and narrow than when the content state equals the mean initial opinion. If the content state is too extreme, it spreads very little.

### v2

@@ V Quantifying content spread
Now that we have predicted the onset of infodemics in our content-spreading model on configuration-model networks, we measure the sizes (i.e., the numbers of nodes) in the resulting dissemination trees. The size of a dissemination tree equals the total number of content shares. Studying dissemination trees allows us to explore the long-term behavior of our content-spreading model. These calculations thereby complement our examination of early-stage spreading in Section IV.
@@ V.1 Analysis of the total number of content shares
Given the number of nodes and the degree distribution of a configuration-model graph $G$, a content state $x_{0}$, a receptiveness parameter $c$, and a distribution $\phi(x)$ of opinions, we again use generating functions (as in Section IV.1) to estimate the total number of content shares (i.e., the size) of a dissemination tree $G_{0}$. This analysis treats content spread as a percolation process. See Ref. [Newman, 2018] for a detailed description of this approach.
Let the “spreading set” $\mathcal{S}$ of a network denote the set of nodes that have shared a piece of content. The spreading set $\mathcal{S}$ has $SN$ elements (i.e., there are a total of $SN$ content shares), where $S$ is the fraction of a network’s nodes that have shared that content. Node opinions are independent of connectivity in our content-spreading model, so the probability that a node is in $\mathcal{S}$ equals the probability that it is receptive to the content multiplied by the probability that it belongs to the connected component of a source node. In this calculation, we assume that there is a single source node. Therefore, $SN$ is the size of the connected component of one source node. Let $u$ be the mean probability of not being connected to this component via a particular neighbor. We then have
@@
where we recall that $p_{k}$ is the probability that a node has degree $k$ and $g_{0}(u)$ is the generating function of the degree distribution of the graph.
To calculate $S$, we need to determine $u$. A particular node $i$ is not connected to the spreading set $\mathcal{S}$ via a particular node $j$ either because it is not receptive to the content or because it is not in the same component as the content. In the latter case, node $i$ is isolated from the content spread because of non-receptive nodes. For a node $j$ with $k$ edges aside from the one to node $i$, the probability that one of these scenarios occurs is $1-s(x_{0},c)+s(x_{0},c)u^{k}$. Therefore,
@@
where $q_{k}$ again denotes the probability of having $k$ neighbors other than the spreading edge (i.e., the excess degree is $k$) and $g_{1}(u)$ again denotes the generating function of the excess degree distribution. By inspection, $u=1$ is always a solution; this solution entails not having an infodemic. We are interested in whether or not there are also solutions $u\in(0,1)$.
In some special cases, it is possible to solve for $u$ analytically, but it is typically difficult (or even impossible) to do so even when $g_{1}(u)$ has a simple closed-form expression. However, one can find a root $u^{*}\in(0,1)$ of Section V.1 numerically either using an explicit expression or using an approximation of the generating function of the excess degree distribution Newman (2018). One then substitutes $u^{*}$ into Equation 11 to obtain an approximation of $S$. The solution $u^{*}\in(0,1)$ yields an approximation of $S$ when there is an infodemic.
In Figure 4, we show an example in which we compare our approximation of the total number of content shares ($SN$) to numerical simulations in configuration-model networks of three different sizes. For each size (200, 2400, and 4800 nodes), we generate 1000 configuration-model networks with degree sequences from a Poisson distribution with mean $\lambda=5$. We plot histograms of the total number of content shares, and we observe good agreement between our analytical and numerical results.
Figure 4: Histograms of the total number of content shares in our content-spreading model across 1000 trials for 200-node, 2400-node, and 4800-node configuration-model networks. The dashed vertical lines indicate the numbers of content shares (i.e., the dissemination-tree sizes) that we predict from our analysis. In each realization, we generate an $N$-node configuration-model network with a degree sequence that we obtain from a Poisson distribution with mean $\lambda=5$. We draw the node opinions $x_{i}$ uniformly at random from $(0,1)$, and we set the content state to $x_{0}=0.5$ and the receptiveness parameter to $c=0.2$. In each realization, we draw a new degree sequence and a new set of node opinions.
@@ VI Summary statistics for dissemination trees
Figure 5: The effect of varying the network size $N$ on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spreading model on configuration-model networks. The solid curves give means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate an $N$-node configuration-model network with a degree sequence from a Poisson distribution with mean $\lambda=5$. We vary $N$ from $100$ to $5000$ in increments of $100$. Each realization has different initial node opinions, which we draw uniformly at random from $(0,1)$. The content state is $x_{0}=0.5$ and the receptiveness parameter is $c=0.2$. In each realization, we draw a new degree sequence and a new set of node opinions. Both the total number of content shares (i.e., the dissemination-tree size) and the width grow linearly with $N$. The longest-path length grows quickly at first as we increase $N$, and then it grows much more slowly with $N$. The structural virality appears to saturate at a constant value for sufficiently large $N$.
@@
Figure 6: The effect of varying the expected mean degree $\lambda$ of a network on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spreading model on configuration-model networks. The solid curves give means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate a 2000-node configuration-model network with a degree sequence from a Poisson distribution with mean $\lambda$. We vary $\lambda$ from $1$ to $40$ in increments of $1$. Each realization has different initial node opinions, which we draw uniformly at random from $(0,1)$. The content state is $x_{0}=0.5$ and the receptiveness parameter is $c=0.2$. In each realization, we draw a new degree sequence and a new set of node opinions. The total number of content shares grows initially and then saturates at about $800$. The width also tends to increase with $\lambda$, with a possible plateau in a small interval near $\lambda=20$. The longest-path length and structural virality increase at first and then decrease, eventually leveling off at a constant value for sufficiently large expected mean degrees. In our simulations, the longest-path length has a maximum at about $\lambda=4$ and the structural virality has a maximum at about $\lambda=3$.
@@
Figure 7: The effect of varying the receptiveness parameter $c$ on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spreading model on configuration-model networks. The solid curves give means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate a 2000-node configuration model network with a degree sequence from a Poisson distribution with mean $\lambda=5$. Each realization has different initial node opinions, which we draw uniformly at random from $(0,1)$. The content state is $x_{0}=0.5$. We vary $c$ from $0.02$ to $0.48$ in increments of $0.02$. In each realization, we draw a new degree sequence and a new set of node opinions. The total number of content shares (i.e., the dissemination-tree size) and the width are very small until about $c=0.1$, and then they grow with $c$. Both growth rates seem roughly linear after some earlier slow growth, although the curve for the total number of content shares seems to be concave down. The longest-path length increases initially with $c$ before reaching a maximum and then decaying. The structural virality also appears to increase initially with $c$ before reaching a maximum and then decaying to a constant value. The standard deviations of the longest-path length and the structural virality are large.
@@
Figure 8: The effect of varying the content state $x_{0}$ on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spreading model on configuration-model networks with initial node opinions from a uniform distribution. The solid curves show means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate a 2000-node configuration-model network with a degree sequence from a Poisson distribution with mean $\lambda=5$. Each realization has different initial node opinions, which we draw uniformly at random from $(0,1)$. The receptiveness parameter is $c=0.2$. We vary $x_{0}$ from $0.02$ to $0.98$ in increments of $0.02$. In each realization, we draw a new degree sequence and a new set of node opinions. When $x_{0}\in(0.2,0.8)$, the total number of content shares, the width, the longest-path length, and the structural virality are all roughly constant. This is not surprising because we draw initial node opinions uniformly at random, so the transmission probability for content to spread along an edge depends only on $c$ in this interval (see Equation 4). The symmetry of these summary statistics with respect to $x_{0}$ is also clear from Equation 4. Content states that are closer to the boundary of opinion space produce dissemination trees with fewer total content shares, smaller widths, and shorter longest paths. However, the structural virality of such “extreme” content states is slightly larger.
@@
Figure 9: The effect of varying the content state $x_{0}$ on the total number of content shares, the width, the longest-path length, and the structural virality of dissemination trees of our content-spreading model on configuration-model networks with initial node opinions that we draw from a Gaussian distribution with mean $0.5$ and standard deviation $0.1$. In this figure, the opinion space is $\mathbb{R}$. This opinion space and the initial opinion distribution differ from the ones in Figure 8. The solid curves give means across $1000$ realizations, and the shaded regions give the standard deviations. In each realization, we generate a 2000-node configuration-model network with a degree sequence from a Poisson distribution with mean $\lambda=5$. The receptiveness parameter is $c=0.2$. We vary the content state $x_{0}$ from $0.02$ to $0.98$ in increments of $0.02$. In each realization, we draw a new degree sequence and a new set of node opinions.
@@
In this section, we perform numerical experiments and study the output of our bounded-confidence content-spreading model on configuration-model networks. We quantify its features by computing four summary statistics of the dissemination trees that we obtain in our simulations.
@@
  • Length of a longest adoption path. An adoption path is a path in a dissemination tree from a source node to another node in the tree Oh and Porter (2018). We measure the “depth” of content spread by calculating the length of a longest adoption path.
  • Width. One can arrange a dissemination tree $G_{0}$ with the source node at the top (i.e., at level 00), nodes with adoption-path length $1$ at level $1$, nodes with adoption-path length $2$ at level $2$, and so on. The width of a dissemination tree is the largest number of nodes that adopt the content in a level. Mathematically, the width is
    $\max_{{b}\in\{1,\ldots,l_{*}\}}\text{(number of nodes with {adoption-path} length}{\,\,b})\,,$
    where $l_{*}$ is the length of a longest adoption path.
  • Structural virality. Structural virality (i.e., the Wiener index) was used by Goel et al. Goel et al. (2016) to measure the viral nature of content spread. The structural virality $v$ is the mean shortest-path length between nodes in a dissemination tree $G_{0}$. Mathematically, the structural virality is
@@
One natural question to ask is how the number of nodes of a network affects the spread of content on that network. We study this question by increasing the network size $N$ for fixed expected mean degree $\lambda$ (see Figure 5). With this construction, the expected number of edges remains constant as we increase $N$. As implied by our analysis in Section V, the total number of content shares grows linearly with $N$ (with a slope indicated by Equation 11). The dissemination-tree width also appears to grow linearly with $N$. The longest-path length and structural virality change more dramatically for small $N$ than for large $N$. The former appears to grow very slowly for large $N$, and the latter appears to eventually saturate at a constant value.
@@
In Figure 6, we examine the impact of increasing the expected mean degree $\lambda$ on the dissemination-tree statistics. The total number of content shares increases with $\lambda$ before saturating at a constant value. The width also tends to increase with $\lambda$, with a possible plateau in a small interval near $\lambda=20$. The longest-path length and structural virality have maxima for small values of $\lambda$. When $\lambda$ is large, it is seemingly possible for the content to be adopted by all receptive and connected nodes in just a few steps, yielding a large dissemination tree that reaches many new nodes in a single time step, short adoption paths, and small structural viralities. The maxima in the longest-path length and structural virality indicate that a spreading process persists longer (and hence travels farther from a source node) when nodes have fewer neighbors on average (but still have a mean degree that is large enough for an infodemic to occur).
@@
Varying the receptiveness parameter $c$ (see Figure 7) yields similar dissemination-tree statistics as varying the expected mean degree. We again observe maxima in the longest-path length and the structural virality. These maxima likely arise because larger receptiveness values favoring faster content spread, although the longest-path length and structural virality both have large standard deviations in this experiment. We observe evidence of a phase transition in the total number of content shares at the value of the receptiveness parameter $c$ that we predicted in our analysis in Section IV.1. There also appears to be an accompanying phase transition in the dissemination-tree width. In these simulations, the content state is $x_{0}=0.5$. We observe qualitatively similar behavior with content states (e.g., $x_{0}=0.15$) that are closer to the boundary of opinion space but still in the infodemic regime. In our repository (see https://github.com/hzinnbrooks/bounded-confidence-spreading-process/supplementary-figures), we include analogous figures to Figure 6 and Figure 7 for such content states.
@@
We also examine the impact of the content state on our dissemination-tree statistics for two different initial opinion distributions (see Equation 3). In Figure 8, we show results of simulations in we draw the initial node opinions from a uniform distribution on the interval $(0,1)$. In Figure 9, we show results of simulations in which we draw the initial node opinions from $\mathbb{R}$ using a Gaussian distribution with mean $0.5$ and standard deviation $0.1$. As expected, we observe smaller values of all four summary statistics when the content state lies in the tails of a opinion distribution $\phi(x)$ (or near the boundary of the distribution interval for the uniform distribution). We observe maxima in the total number of content shares and dissemination-tree width at the distribution mean (and also for a very large interval around it for the uniform distribution). For the Gaussian distribution, the content states that yield the largest longest-path lengths and structural viralities occur away from the mean (and are symmetric), indicating that content can spread farther from a source node in these situations then when the content state equals the mean initial opinion. These dissemination trees are longer and narrow than when the content state equals the mean initial opinion. If the content state is too extreme, it spreads very little.